OpenStudy (anonymous):

To approach a runway, a plane must begin a 7 degree descent starting from a height of 3 miles above the ground. To the nearest mile, how many miles from the runway is the airplane a the start of this approach?

OpenStudy (anonymous):



OpenStudy (anonymous):

h/x=sin7 x=h/sin7=3/sin7